# The Necessity of Imperfection: Reversing Model Collapse via Simulating Cognitive Boundedness

### v2

@@ 6.5 Ethics and Safety
The PMCSF framework proposed in this study, while innovative, constitutes a double-edged sword. The high-fidelity text generated by the CTE module, which possesses strong anti-AI-detection capabilities, could, if misused, facilitate large-scale disinformation campaigns—an risk exacerbated by the framework’s ability to mimic human cognitive texture. Furthermore, as a paradigm that requires no model fine-tuning, the PMCSF framework has a relatively low barrier to replication; malicious actors could exploit its market sentiment analysis and prediction capabilities for market manipulation or to gain unfair trading advantages, posing threats to financial security.
In light of these risks and out of consideration for AI safety and responsible release, this paper has chosen not to disclose the full original prompt text for the CTE module. However, detailed algorithmic pseudo-code and architectural logic are provided in Appendix B; these materials are sufficient to support academic replication and verification while minimizing the risk of malicious misuse. We call upon the academic community to establish stricter access controls and ethical review mechanisms when developing such powerful generative tools—measures that are essential for preventing abuse while preserving the framework’s scientific value.
@@ C.1 Calibration Methodology & Data Inventory
  • GARCH Baseline Parameters: A hybrid calibration method was deployed, whereby GJR-GARCH(1,1) was directly estimated via Maximum Likelihood Estimation (MLE) for data-rich typical quadrants (e.g., the 2015 stock market crash), while behavioral finance theories (e.g., loss aversion coefficients) guided expert calibration for data-sparse quadrants.
@@
C.2 GJR-GARCH Parameter Calibration for Core Dimensions
Table 5 presents the estimated GJR-GARCH parameters for the Fear dimension across distinct market quadrants—parameters that were not only calibrated for typical quadrants (e.g., Full Bubble or Structural Tear) but also interpolated and theoretically derived for all six quadrants to enable all-weather market simulation. This approach ensures the framework can model both common and extreme sentiment states, a critical feature for avoiding the “statistical mode collapse” discussed in the main text.
Table 5: GJR-GARCH Parameter Calibration Results for Fear Dimension (Full Version)
Macro-Quadrant Dominant Emotion $\omega$ $\alpha$ $\alpha^{-}$ $\beta$ Calibration Basis
A. Full Bubble Greed 0.03 0.05 0.15 0.88 2021 Beta + Expert Setting
B. Structural Tear Joy 0.01 0.10 0.08 0.80 2021 Fitting ($\omega,\alpha^{-},\beta$)
C. Dead Ice Fear 0.02 0.05 0.00 0.85 2018 Fitting $\to$ GARCH(1,1)
D. Inert Recession Sadness 0.04 0.03 0.10 0.92 2015 Beta + Expert Setting
E. Recessional Tear Fear 0.05 0.02 0.20 0.90 2015 Anchor (Expert Calib.)
F. Structural Rally Joy 0.01 0.12 0.07 0.78 2021 Fitting ($\omega,\alpha,\alpha^{-},\beta$)
Statistical Insight: In bearish quadrants—notably Recessional Tear (E) and Dead Ice (C)—the asymmetric shock coefficient $\alpha^{-}$ exceeds the symmetric coefficient $\alpha$ at a statistically significant level ($p<0.01$), mathematically lending credence to the asymmetric dynamic wherein panic propagates more rapidly than recovery in bear markets. This finding aligns with behavioral finance theories positing amplified loss aversion during market downturns, a heuristic explicitly modeled in the PMCSF framework.
C.3 Calibration of Satellite Sentiment Dynamics Coefficients
@@
C.4 Holiday Effects and Decay Coefficient Calibration
@@
C.4.1 Post-Holiday Volatility Effect
@@
C.5 Shock Response Model Parameter Calibration

### v3

@@ Abstract
Abstract
The preceding Appendices A-G provided the theoretical definitions, mathematical formalizations, and qualitative context (e.g., the Lexicon of Irrationality) essential for understanding the financial and cognitive logic of this study.
The following Supplementary Materials (Version 2.0) focus on the system’s engineering implementation. This section includes algorithmic pseudocode, system architecture diagrams, and granular parameter calibration logs, intended for computer scientists and quantitative developers seeking technical reproducibility.
While the following materials provide the comprehensive architectural logic required for scientific reproducibility, certain executable components have been deliberately obfuscated or withheld.
Specifically, the "Financial Sentiment Injection Protocol" and the "Micro-Chaos Lexicon (v.Alpha)" detailed in Section S3 have demonstrated a capability to induce self-reinforcing panic loops in high-frequency algorithmic trading environments during internal stress tests.
To prevent the potential weaponization of this framework as a "Cognitive Dirty Bomb" for market manipulation, the raw prompt sequences in this document have been sanitized. The provided pseudocode is sufficient for academic verification but has been inerted for deployment in live trading systems.
Figure 1: The Topology of Survival: A Visual Abstract of “Cognitive Phase Transition” During Financial Crisis. This 3D manifold illustrates the core thesis of this study: in non-ergodic markets (e.g., the 2015 Crash), survival is a topological bifurcation. The Purple Zone represents the "Consensus Phase" where both agents behave similarly. However, as the system approaches a critical singularity (June 15):
(A) The Rationality Trap (Red Trajectory): The Human-Baseline strategy, which operates under the constraint of “mean reversion bias” (a heuristic detailed in the study’s glossary), misclassifies structural collapse as transient market noise—a failure to distinguish between cyclical fluctuations and irreversible decline. This misperception propels the strategy toward a gravitational well of collapse, where the pull of declining asset values becomes inescapable (drawdown: -23.2%).
(B) The Instinctual Escape (Blue Trajectory): The CTE-Enhanced strategy, infused with synthetic cognitive noise (a key component of the PMCSF framework), initiates a phase transition upon detecting a GARCH volatility peak ($h_{joy}\to 0.92$)—a threshold where the model’s “cognitive texture” (imperfections mimicking human heuristic biases) triggers a break from statistical optimality. Shattering the symmetry of mean-reverting expectations, the strategy implements a “Digital Spartan” escape: full liquidation that avoids the gravitational well’s pull.
Metaphorical Insight: Survival is not about better fitting; it is about breaking the loop.
@@ 1.4 Contributions of This Study
  • Open Source Contribution: To facilitate reproducibility and provide an intuitive understanding of the proposed Cognitive Texture mechanism, we release Adversarial-Text-Protocols, an open-source prompt engineering suite. This allows researchers to interactively experience the effects of micro-perturbations on LLM outputs via standard interfaces.
@@ 6.2 Implications for AI Development: From “Probability Statistics” to “Cognitive Dynamics”
6.3 Ecological Validity: A Computational Echo of Adaptive Markets
Initially designed to address the data collapse problem afflicting generative AI, this study’s experimental results nonetheless provide unique computational evidence for Andrew Lo’s Adaptive Market Hypothesis (AMH)—a theory positing that the irrationality of market participants is an evolutionarily adaptive trait[25]. Our CTE-Enhanced strategy—detailed in Figure 7—corroborates this: it was precisely the seemingly “imperfect” cognitive noise (synthetic panic) that activated the GARCH model’s stress response mechanism, allowing the algorithm to survive the stock market crash. This implies the PMCSF framework has inadvertently built a bridge between AI’s “model robustness” and finance’s “market survival theory.”
Figure 7: Real-time Parameter Drift (Fear Dimension). Visualizing the “cognitive phase transition” within the CTE-Enhanced strategy. The blue line ($\beta$) shows memory persistence shifting to a PTSD-like state, while the red dashed line ($\alpha^{-}$) quantifies elevated loss aversion during the crisis.
Figure 7 visualizes the real-time parameter drift of the fear dimension, revealing a “cognitive phase transition” within the CTE-Enhanced strategy:
  • Variation of Memory Persistence ($\beta$, Blue Line): At the June 15th crisis onset, the memory parameter $\beta_{fear}$ displayed sharp fluctuations—an indicator of the system’s recognition of a structural environmental break, switching it from a “stationary mode” to a “crisis mode.” Subsequently, the $\beta$ value stabilized at a high level ($0.87+$), indicating the agent entered into a “long-memory (PTSD-like)” state—one retaining an extremely long half-life for past fear signals and thus immune to subsequent bull traps.
  • Quantification of Loss Aversion ($\alpha^{-}$, Red Line): The red dashed line delineates that the asymmetric shock coefficient $\alpha^{-}_{fear}$ remained consistently elevated ($0.10\text{--}0.11$) throughout the crisis period. Compared to the baseline bubble-phase value (typically $\approx 0.05$), this parameter doubled—meaning the agent assigned to “bad news” a volatility weight more than twice that of “good news” in crisis mode. This constitutes the mathematical expression of a survival discipline: “better to sell mistakenly than to catch a falling knife mistakenly.”
@@ 6.6 Ethics and Safety
While the PMCSF framework introduces significant methodological advancements, it also presents a potential dual-use risk. The high-fidelity text generation capabilities of the CTE module, characterized by robust resistance to AI detection, could theoretically be weaponized for large-scale disinformation campaigns. Moreover, given that the framework operates without the need for model fine-tuning, its low deployment barrier raises concerns that malicious actors could exploit its sentiment analysis capabilities for market manipulation or to gain asymmetric trading advantages, thereby threatening financial stability.
Adhering to principles of responsible AI disclosure, we have elected to redact the full original prompt text of the CTE module. To further mitigate risks, the open-source release is restricted to a "General Narrative Mode." Sensitive components used in our A-share stress tests—specifically the "Financial Sentiment Injection Protocol" and the "Micro-Chaos Lexicon"—have been withheld from the public repository to prevent their weaponization in high-frequency trading contexts.
However, to ensure scientific reproducibility, we provide comprehensive algorithmic pseudocode and architectural logic in Appendix B and the Supplementary Materials. We believe this strikes an optimal balance between enabling academic verification and preventing malicious misuse. We explicitly call for the establishment of rigorous access controls and ethical review standards for similar generative frameworks, ensuring that scientific progress does not come at the cost of safety.
@@ 7.4 Concluding Remarks
Furthermore, this study uncovers a profound convergence between artificial and biological intelligence: the survival necessity of imperfection. This finding resonates with the evolutionary perspective of financial markets, suggesting the path to Artificial General Intelligence (AGI) may lie not in perfect logic but in adaptive boundedness.
@@ C.1 Calibration Methodology & Data Inventory
C.2 Calibration of Satellite Sentiment Dynamics Coefficients
@@
C.3 Holiday Effects and Decay Coefficient Calibration
@@
C.3.1 Post-Holiday Volatility Effect
@@
C.4 Shock Response Model Parameter Calibration